Complete the table for the equation.
5x + y = -20
_________
_X_|_Y__
-5_|_H__
_0_|_H___
_H__|-25
I have place an H where the table needs completed. There are three of them.

Answer by Nate(3500)

5x + y = -20
_X_|_Y__
-5_|_H__
5(-5) + y = -20
-25 + y = -20
y = 5
_X_|_Y__
-5_|_5__
_X_|_Y__
_0_|_H__
5x + y = -20
5(0) + y = -20
y = -20
_X_|__Y__
_0_|_-20_
_X_|__Y__
_H_|_-25_
5x + y = -20
5x + -25 = -20
5x = 5
x = 1
_X_|__Y__
_1_|_-25_